What is the concentration in moles per liter of a solution that contains 7.68 g of H2SO4 in 150.0 mL of water?


1
Expert's answer
2021-05-27T07:34:33-0400
  • Step1: Find the molar mass of H2SO4
  • Atomic mass of H=1, S=32 and O=16
  • Molar mass= 2(1)+32+4(16)=98g/mol
  • Step 2: Convert 7.68g to moles
  • If 98g are in 1 mole, what about 7.68g
  • 7.68x1/98=0.0784 moles
  • Step 3: Find the concentration in moles per liter.
  • If 0.0784 moles are in 150ml, what about 1000ml (1 liter).
  • 0.0784 x1000/150=0.5224moles/liter